OpenStudy (anonymous):

A feather is dropped onto the surface of the moon. How far will the feather have fallen if it reaches the surface in 9.00 seconds?

OpenStudy (anonymous):

you must use the equations of kinematics, in particular the uniformly accelerated motion equation, but you must know the acceleration of gravity on the moon...

OpenStudy (anonymous):

s = (1/2) a t^2 a = Earth/6 = (9.8/6) m/s^2

OpenStudy (gebooors):

To be exact, in Moon g = 1,62 m/s^2